# Perceptions of the Fairness Impacts of Multiplicity in Machine Learning

### v1

@@ B.1 Tasks
B.2 Part 1: Fairness questions
@@
B.3 Part 1: Comprehension questions
We ask one comprehension question at the beginning of the survey. Participants who do not answer correctly after two tries are not allowed to complete the survey. The question is, “What does model multiplicity refer to?” with these choices: (a) It refers to a model making multiple correct predictions, (b) It means there are several models that perform equally well but makes different mistakes, (c) It describes an AI model that performs better than all other models, and (d) It refers to an AI model that can perform multiple task simultaneously. The correct answer is (b), which can be found in the final paragraph of the text provided to the user, almost word-for-word.
We also ask two critical-thinking questions midway through Part 1 of the survey to encourage participants to slow down and better absorb the material. The first question is multiple-choice, asking which of four statements is true: (a) The original AI model is the only good way to do [task], (b) The new AI and human experts always choose the same people, (c) Multiple models are equally good at [task], or (d) The original AI and new AI never agree on [task]. The correct answer is (c) and this is visible to participants as soon as they make any selection. The open-ended question asks “If the original AI and the new AI disagree about whether or not to [predict + or -], how should the [decision-making entity] decide what to do?” Respondents much enter at least 8 characters in their answer, but apart from that requirement there is no in-survey validation of the answer.
B.4 Part 2: Costs
@@ Appendix C Alternate statistical analyses
We describe two alternate statistical analyses for the Part 1 data (change in fairness perceptions after learning about multiplicity). As stated in section 4.4, we preregistered our study with a plan to use linear mixed models. Those results are in the main text. We include two alternate analyses here for two reasons: first, there is disagreement in the statistical community about what type of analysis is appropriate for Likert-scale data [9, 14, 46, 54, 70]. Including multiple options serves to “future-proof” our analysis in the case of future statistical consensus. Second, while preregistration is one tool to guard against “researcher degrees of freedom” [89], another option is a multiverse analysis [92]. In a true multiverse analysis, researchers present the outcomes of all feasible statistical analyses on their data. Then, if the same conclusions hold for all or most of the analyses, they can be considered more robust. We present a partial multiverse analysis.
C.1 ART-ANOVA
Aligned-Rank-Transform (ART) ANOVA [98] is a procedure that transforms ordinal (e.g., Likert-scale) data to fit the ANOVA assumptions. We fit our ART-ANOVA model and perform post-hoc contrast tests [30] using the ARTool package in R [49]. Our formula includes the factors time, task stakes (binary), task uncertainty (binary), and task framing and all possible interactions between factors. We also include the participant ID to account for within-subjects factor. We test H using a post-hoc contrast test with “time” as the factor of interest. We test H, H, and H with a post-hoc contrast test of the interaction term between “time” and the relevant task characteristic (with the “Interaction” flag set to true).
@@ Results
C.2 Non-parametric tests
The Mann-Whitney U-test [64] is another way we can evaluate H-H. This test is non-parametric, so does not require any assumptions about the data’s normality or distribution. We perform the tests using the wilcox.test function in R. For H, we test whether the average fairness perception is different as stratified by “time”. For H-H, we compute the change in score over time for each sample, and then conduct a Mann-Whitney U test on the data stratified by the relevant task characteristic.
@@ Appendix D Additional results (Fairness perceptions)
D.1 Task choice and baseline fairness perceptions
@@ E.2 Qualitative results
| Care | Job | Tumor | Tax | Upgrade | Rebook | App | Reviews | Total | |
| plan | probation | detection | audit | ||||||
@@
E.3 Task-by-task preferences
@@
E.4 Costs
As noted in finding E4 in section 5.3, no extra cost is significantly preferred over extra cost in aggregate. However, for three high-stakes settings (tumor detection, care plan access, and job probation), this distinction is not significnat. The point estimates for marginal mean preferences for extra cost are 0.436 (SE 0.032, p¡0.0001) for tumor detection, 0.474 (SE 0.038, p¡0.0001) for care plan access, and 0.445 (SE 0.035, p¡0.0001) for job probation. (A point estimate less than 0.50 indicates that no extra cost is more preferred, and a point estimate of 0.50 indicates that there is no preferences between the cost options.)

### v2

@@ B.1. Tasks
B.2. Multiplicity education
We provide the following text to participants at the beginning of the survey:
The image shows a table with 5 columns and two rows. The column headers are pictures, each of a cat or a dog. The two rows are labels AI#1 and AI#2. The table cells indicate what the AI models predict as the label for each image. Each cell includes the prediction (i.e., the word “cat” or the word “dog”) as well as a green checkmark, in the case of a correct prediction, or a red X in the case of an incorrect prediction. AI#1 is correct on 4/5 examples (it identifies one dog (column 3) as a cat). AI #2 is also correct on 4/5 examples, but identifies one cat (column 2) as a dog. Text to the right of each row points out that both AIs are correct on 4/5 examples.
When building Artificial Intelligence (AI), choosing the best model is crucial. AI developers consider a range of AI models and choose the one with the best performance. For example, if the goal is to distinguish between cats and dogs, developers will test different AIs using a set of images and choose the AI that correctly identifies the most pictures.
In the example below, AI #1 correctly identifies 4 out of 5 images. Suppose this is the highest-performing AI model available.
We display the top half of fig. 4 (i.e., images and AI #1 only).
Sometimes, however, multiple AI models perform equally well but differ in their specific errors. This is called model multiplicity. For example, AI #2 also correctly identifies 4 out of 5 images but makes a different mistake than AI #1.
We display fig. 4.
AI #1 and AI #2 are both well-suited for classifying cat and dog images. If no other AI performs better, developers must choose between AI #1 and AI #2.
Model multiplicity is common in AI: multiple models can perform equally well overall but make different mistakes. In some settings, like predicting cats and dogs, model multiplicity is inevitable but choosing between models is low-stakes. In other settings, choosing between models is more challenging, such as when the model predictions directly affect people.
Then, we ask the comprehension question “What does model multiplicity refer to?” with these choices: (a) It refers to a model making multiple correct predictions, (b) It means there are several models that perform equally well but makes different mistakes, (c) It describes an AI model that performs better than all other models, and (d) It refers to an AI model that can perform multiple task simultaneously.
B.3. Part 1: Fairness questions
@@
B.4. Part 1: Critical-thinking questions
We ask two critical-thinking questions midway through Part 1 of the survey to encourage participants to slow down and better absorb the material. The first question is multiple-choice, asking which of four statements is true: (a) The original AI model is the only good way to do [task], (b) The new AI and human experts always choose the same people, (c) Multiple models are equally good at [task], or (d) The original AI and new AI never agree on [task]. The correct answer is (c) and this is visible to participants as soon as they make any selection. The open-ended question asks “If the original AI and the new AI disagree about whether or not to [predict + or -], how should the [decision-making entity] decide what to do?” Respondents much enter at least 8 characters in their answer, but apart from that requirement there is no in-survey validation of the answer.
B.5. Part 2: Costs
@@ Appendix C Alternate statistical analyses
We describe two alternate statistical analyses for the Part 1 data (change in fairness perceptions after learning about multiplicity). As stated in section 4.4, we preregistered our study with a plan to use linear mixed models. Those results are in the main text. We include two alternate analyses here for two reasons: first, there is disagreement in the statistical community about what type of analysis is appropriate for Likert-scale data (Blanca Mena et al., 2017; Carifio and Perla, 2008; Jamieson, 2004; Kuzon et al., 1996; Norman, 2010). Including multiple options serves to “future-proof” our analysis in the case of future statistical consensus. Second, while preregistration is one tool to guard against “researcher degrees of freedom” (Simmons et al., 2011), another option is a multiverse analysis (Steegen et al., 2016). In a true multiverse analysis, researchers present the outcomes of all feasible statistical analyses on their data. Then, if the same conclusions hold for all or most of the analyses, they can be considered more robust. We present a partial multiverse analysis.
C.1. ART-ANOVA
Aligned-Rank-Transform (ART) ANOVA (Wobbrock et al., 2011) is a procedure that transforms ordinal (e.g., Likert-scale) data to fit the ANOVA assumptions. We fit our ART-ANOVA model and perform post-hoc contrast tests (Elkin et al., 2021) using the ARTool package in R (Kay et al., 2021). Our formula includes the factors time, task stakes (binary), task uncertainty (binary), and task framing and all possible interactions between factors. We also include the participant ID to account for within-subjects factor. We test H using a post-hoc contrast test with “time” as the factor of interest. We test H, H, and H with a post-hoc contrast test of the interaction term between “time” and the relevant task characteristic (with the “Interaction” flag set to true).
@@ Results
C.2. Non-parametric tests
The Mann-Whitney U-test (Mann and Whitney, 1947) is another way we can evaluate H-H. This test is non-parametric, so does not require any assumptions about the data’s normality or distribution. We perform the tests using the wilcox.test function in R. For H, we test whether the average fairness perception is different as stratified by “time”. For H-H, we compute the change in score over time for each sample, and then conduct a Mann-Whitney U test on the data stratified by the relevant task characteristic.
@@ Appendix D Additional results (Fairness perceptions)
D.1. Task choice and baseline fairness perceptions
@@
D.2. Demographic factors
We consider binary gender, integer age, binary experience with AI (defined as whether the participant currently uses AI in their work or school), and binary education (whether the participant has a bachelor’s degree or higher). There is no significant difference in initial fairness perceptions based on gender (t-test p=0.822), age (linear regression p=0.184), experience with AI (t-test p=0.088), or education (t-test p=0.328). There is no significant different in change in fairness scores after learning about multiplicity for gender (t-test p=0.921), age (linear regression p=0.944), AI experience (t-test p=0.460), or education (t-test p=0.413).
@@ E.2. Qualitative results
| Care plan | Job | Tumor | Tax audit | Upgrade | Rebook | App | Reviews | Total | |
| probation | detection | ||||||||
@@
E.3. Task-by-task preferences
@@
Marginal mean preferences over multiplicity resolution techniques stratified by task stakes, task framing, and task uncertainty. We report which results have non-overlapping confidence intervals for each setting. For task stakes, human (more preferred in high stakes), default bad (more preferred in low stakes), zero cost (more preferred in low stakes) and non-zero cost (more preferred in high stakes). For task framing, random (more preferred for reward tasks), default good (more preferred for reward), and default bad (more preferred for punishment). For task uncertainty, no factor has non-overlapping confidence intervals.
E.4. Costs
As noted in finding E4 in section 5.3, no extra cost is significantly preferred over extra cost in aggregate. However, for three high-stakes settings (tumor detection, care plan access, and job probation), this distinction is not significant. The point estimates for marginal mean preferences for extra cost are 0.436 (SE 0.032, p¡0.0001) for tumor detection, 0.474 (SE 0.038, p¡0.0001) for care plan access, and 0.445 (SE 0.035, p¡0.0001) for job probation. (A point estimate less than 0.50 indicates that no extra cost is more preferred, and a point estimate of 0.50 indicates that there is no preferences between the cost options.)
E.5. Demographic factors
There is no significant difference in preference for multiplicity resolution technique based on education level or AI experience. That is, a t-test comparing the results for each resolution technique has a p-value greater than 0.05. When we consider age as a binary factor (divided at 35, roughly the median age and a common cutoff for “young adult”), fancy is preferred by people younger than 35 (diff=0.081, SE=0.039, t= 2.054, p=0.040) and default_good is preferred by people older than 35 (diff=-0.101, SE=0.040, t=-2.515, p=0.012). Looking at gender, fancy is preferred by men (diff=0.0783, SE=0.039, t=2.009, p=0.045) and random is preferred by women (diff=-0.070, SE=0.035, t=-2.027, p=0.043). We encourage future work to explore whether these trends hold as the experiment setup is varied.